### v1

@@ 4.2 Ca2Y2Cu5O10
Nevertheless, a notable point is that $J_{2}/\lvert J_{1}\rvert=0.23$ is very close to the critical point of $0.25$. Therefore, tuning $J_{2}/\lvert J_{1}\rvert$ to exceed $0.25$ may not be unrealistic. For instance, applying pressure might achieve this. Within the effective spin Hamiltonian, the smaller couplings are expected to be most sensitive to pressure. Consequently, it can be anticipated that the change in $J_{2}$ would be more significant than that in $J_{1}$ under pressure, making it feasible to achieve $J_{2}/\lvert J_{1}\rvert>0.25$. Once $J_{2}/\lvert J_{1}\rvert>0.25$ is realized, the CAFO could be retained by the relatively large interchain couplings $J_{\rm a}+J_{\rm a}\sim 2.29$ meV, potentially leading to the observation of MBS with a relatively large binding energy. Furthermore, if the value of $J_{2}/\lvert J_{1}\rvert$ can be continuously varied by pressure, it might be possible to observe MBS ground state that successively changes $p$ with pressure.
@@ 4.3 LiCuSbO4
The third material under consideration is LiCuSbO4 [59]. In this compound, the magnetic couplings between layers are negligible, allowing it to be effectively treated as a 2D system. The structure of the interchain couplings is equivalent to that depicted in Fig. 1e. The intrachain couplings have been estimated as $J_{1}=-125$ K and $J_{2}=37.6$[23], leading to a ratio of $J_{2}/\lvert J_{1}\rvert=0.3$. Given this ratio is very close to the critical value of 0.25, even moderately strong AFM interchain couplings could stabilise a CAFO state. However, in practice, LiCuSbO4 exhibits short-range incommensurate spin correlations below $T\sim 9$ K and, unlike the spiral spin-chain compounds LiCuVO4 [60, 61] and LiCuZrO4 [62, 63], does not show long-range magnetic order down to $T\sim 0.1$ K. This may suggest that the interchain couplings in LiCuSbO4 are very weak. Indeed, DFT and DFT+$U$ band structure calculations estimate $J_{\rm d}\sim 1$[23].
@@
For $J_{\rm d}=0.75$ K, CAFO is not yet stabilised, and $G_{1}(\omega)$ exhibits the lowest energy excitation, indicating the absence of magnon binding. Nonetheless, the proximity of the lowest energy excitation in $G_{3}(\omega)$ suggests that a slight increase in $J_{\rm d}$ might stabilise a triatic ground state. Then, examining the results for $J_{\rm d}=1.25$ K, it is evident that $G_{3}(\omega)$ exhibits the lowest energy excitation, with a relatively large magnon binding energy of approximately 1 meV. Given that LiCuSbO4 is already near the CAFO phase with $J_{2}/\lvert J_{1}\rvert=0.3$, only a modest application of pressure might be required to observe both CAFO and MBS. Interestingly, as $J_{\rm d}$ increases, $G_{1}(\omega)$ shifts to higher energies while changing shape, whereas the spectra for $p>1$ remain largely unchanged. This observation supports the statement that the value of $p$ for the ground state generally corresponds to the number of magnon bindings stabilised by an external field in a single $J_{1}$-$J_{2}$ chain for a given $J_{2}/\lvert J_{1}\rvert$.
@@ 4.4 PbCuSO4(OH)2
The forth material we discuss is PbCuSO4(OH)2, commonly known as linarite [65], which has a 2D layer lattice structure with interchain coupling $J_{\rm d}$, as illustrated in Fig. 1e. The in-chain couplings for this material have been estimated as $J_{1}=-100$ K and $J_{2}=36$[66], with a ratio of $J_{2}/\lvert J_{1}\rvert=0.36$, which is greater than the FM critical point of 0.25. Unlike LiCuSbO4, linarite exhibits long-range spiral order with a small propagation number $q=0.186\pi$ along the $J_{1}$-$J_{2}$ chain below $T_{\rm N}=2.8$ K. This long-range order is believed to be stabilized by a sizable AFM interchain coupling, estimated to be $J_{\rm d}\sim 6$[67]. However, for this material to exhibit CAFO, an interchain coupling of approximately $J_{\rm d}=0.1\lvert J_{1}\rvert\sim 10$ K is required. Similar to LiCuSbO4, there is potential for pressure effects to enhance the interchain coupling. If $J_{\rm d}>10$ K can be achieved, we can explore the possibility of realizing an MBS ground state.
Given the highly anisotropic nature of PbCuSO4(OH)2, we set $\Delta_{1}=0.9$, though the actual anisotropy may be slightly larger. Fig. 5c plots the MES $G_{p}(\omega)$ obtained from DDMRG calculations on a $24\times 4$-chain cluster with $J_{\rm d}=15$ K. The lowest energy excitation state near $\omega=1.3$ meV is given by $G_{2}(\omega)$, indicating that the ground state is a nematic state. This observation aligns with the realization of a nematic state in an external magnetic field for $J_{2}/\lvert J_{1}\rvert=0.36$.
@@ 4.5 Other materials
Until now, we have considered systems where $J_{1}$-$J_{2}$ chains with a bipartite structure transition to CAFO under AFM interchain coupling. However, a simple bipartite structure is not that necessary. For instance, the Q1D mineral antlerite Cu3SO4(OH)4. The magnetic model of this system can be represented as three-leg zigzag ladders, where the central leg hosts competing AFM exchanges, and each side leg features alternating NN FM exchanges $J_{1}$, $J_{1}^{\prime}$ and a weaker NNN AFM exchange $J_{2}$. These three chains are coupled by zigzag-shaped AFM interchain couplings. The side chains are dimerised FM-AFM $J_{1}$-$J_{2}$ chains, with parameters estimated as $J_{1}=-26$, $J_{1}^{\prime}=-11$, and $J_{2}=-6$ (in meV) by GGA+U approximation. Without interchain coupling, the side chains would be in a spiral state [72, 73]. However, the experimentally observed low-temperature magnetic structure shows that the side legs are ferromagnetically polarised in opposite directions, while the central chain is in an AFM order. This suggests that the effect of interchain coupling causes the side chains to polarise ferromagnetically.
@@ 5 Conclusion
In this study, we have demonstrated a novel mechanism for the stabilization of MBS in Q1D cuprates through interchain interactions. Unlike conventional MBS that require an external magnetic field, our results show that AFM interchain coupling can induce MBS even in the absence of an external field. This finding is significant as it provides a new perspective on the interplay between interchain and intrachain interactions in low-dimensional magnetic systems.
Our numerical simulations using the DMRG method confirm that the AFM interchain coupling can act similarly to an internal magnetic field, leading to the formation of CAFO and the subsequent stabilization of MBS. This is particularly intriguing because the AFM interchain coupling has a dual role; while it destabilizes MBS at high fields, it stabilizes them at zero field.
Notably, we found that MBS can be stabilized in Li2CuO2 under ambient pressure and zero magnetic field, highlighting the robustness of this mechanism. However, this aspect has not been experimentally investigated, as it has not been viewed from this perspective before. Additionally, we explored the conditions under which this phenomenon occurs in other materials such as Ca2Y2Cu5O10, LiCuSbO4, and PbCuSO4(OH)2. Our findings suggest that the application of pressure or chemical substitution could further enhance the interchain coupling in these materials, making it feasible to experimentally observe the MBS predicted by our theoretical models.
@@ Methods
We employ the density-matrix renormalization group (DMRG) technique [74] to study the ground-state properties. For single-chain calculations, we study periodic chains with lengths up to $l_{x}=192$ keeping $\chi=800$ density-matrix eigenstates for renormalization procedure. The discarded weight is negligible. When the magnon binding energy is calculated, it is better to use periodic chains to avoid picking up low excitations that may occur near the edges of open chains. For two-chain calculations, we study open ladders with lengths up to $l_{x}=280$ keeping $\chi=3600$. The largest discarded weight is $w_{\rm d}\sim 10^{-11}$. To explicitly fix the direction of magnetization, which corresponds to the direction in which spin-rotation symmetry is broken in each chain, magnetic fields are applied to the sites at both ends of an open ladder. The direction of the magnetic fields is parallel to the z-axis, with fields at both ends of a single chain aligned in the same direction, and those on adjacent chains in opposite directions. However, the magnitude of magnetization extrapolated to the thermodynamic limit is independent of the values of these edge fields. For $12\times 8$-chain calculations, we keep $\chi=1200$ leading to $w_{\rm d}\sim 10^{-10}$. In this case, we focus on the collinear antiferromagnetic ordered state, where the quantum fluctuations are significantly suppressed, the DMRG accuracy is very good. Similar to the case of the open ladder, magnetic fields are applied to both ends of the system.
@@
We employ the dynamical density-matrix renormalization group (DDMRG) technique [75] to calculate the magnon excitation spectrum (MES), defined as We calculate the spin excitation spectrum
@@ Acknowledgments
We thank Ulrike Nitzsche for technical assistance. We thank S.D.Wilson and R.O.Kuzian for fruitful discussion. This project is funded by the German Research Foundation (DFG) via the projects A05 of the Collaborative Research Center SFB 1143 (project-id 247310070).
Authors’ contributions C.E.A. and S.N. performed numerical calculations and analyzed the results. S.N. designed the study and drafted the manuscript in discussion with C.E.A. and S.-L.D.. All authors approved the final manuscript.

### v3

@@ 4.3 LiCuSbO4
The third material under consideration is LiCuSbO4 [60]. In this compound, the magnetic couplings between layers are negligible, allowing it to be effectively treated as a 2D system. The structure of the interchain couplings is equivalent to that depicted in Fig. 1e. The intrachain couplings have been estimated as $J_{1}=-125$ K and $J_{2}=37.6$[23], leading to a ratio of $J_{2}/\lvert J_{1}\rvert=0.3$. Given this ratio is very close to the critical value of 0.25, even moderately strong AFM interchain couplings could stabilize a CAFO state. However, in practice, LiCuSbO4 exhibits short-range incommensurate spin correlations below $T\sim 9$ K and, unlike the spiral spin-chain compounds LiCuVO4 [61, 62] and LiCuZrO4 [63, 64], does not show long-range magnetic order down to $T\sim 0.1$ K. This may suggest that the interchain couplings in LiCuSbO4 are very weak. Indeed, DFT and DFT+$U$ band structure calculations estimate $J_{\rm d}\sim 1$[23].
@@
For $J_{\rm d}=0.75$ K, CAFO is not yet stabilised, and $G_{1}(\omega)$ exhibits the lowest energy excitation, indicating the absence of magnon binding. Nonetheless, the proximity of the lowest energy excitation in $G_{3}(\omega)$ suggests that a slight increase in $J_{\rm d}$ might stabilize a triatic ground state. Then, examining the results for $J_{\rm d}=1.25$ K, it is evident that $G_{3}(\omega)$ exhibits the lowest energy excitation, with a relatively large magnon binding energy of approximately 1 meV. Given that LiCuSbO4 is already near the CAFO phase with $J_{2}/\lvert J_{1}\rvert=0.3$, only a modest application of pressure might be required to observe both CAFO and MBS. Interestingly, as $J_{\rm d}$ increases, $G_{1}(\omega)$ shifts to higher energies while changing shape, whereas the spectra for $p>1$ remain largely unchanged. This observation supports the statement that the value of $p$ for the ground state generally corresponds to the number of magnon bindings stabilised by an external field in a single $J_{1}$-$J_{2}$ chain for a given $J_{2}/\lvert J_{1}\rvert$.
@@ 4.4 PbCuSO4(OH)2
The forth material we discuss is PbCuSO4(OH)2, commonly known as linarite [66], which has a 2D layer lattice structure with interchain coupling $J_{\rm d}$, as illustrated in Fig. 1e. The in-chain couplings for this material have been estimated as $J_{1}=-100$ K and $J_{2}=36$[67], with a ratio of $J_{2}/\lvert J_{1}\rvert=0.36$, which is greater than the FM critical point of 0.25. Unlike LiCuSbO4, linarite exhibits long-range spiral order with a small propagation number $q=0.186\pi$ along the $J_{1}$-$J_{2}$ chain below $T_{\rm N}=2.8$ K. This long-range order is believed to be stabilised by a sizable AFM interchain coupling, estimated to be $J_{\rm d}\sim 6$[68]. However, for this material to exhibit CAFO, an interchain coupling of approximately $J_{\rm d}=0.1\lvert J_{1}\rvert\sim 10$ K is required. Similar to LiCuSbO4, there is potential for pressure effects to enhance the interchain coupling. If $J_{\rm d}>10$ K can be achieved, we can explore the possibility of realising an MBS ground state.
Given the highly anisotropic nature of PbCuSO4(OH)2, we set $\Delta_{1}=0.9$, though the actual anisotropy may be slightly larger. Fig. 5d plots the MES $G_{p}(\omega)$ obtained from DDMRG calculations on a $24\times 4$-chain cluster with $J_{\rm d}=15$ K. The lowest energy excitation state near $\omega=1.3$ meV is given by $G_{2}(\omega)$, indicating that the ground state is a nematic state. This observation aligns with the realisation of a nematic state in an external magnetic field for $J_{2}/\lvert J_{1}\rvert=0.36$.
@@ 4.5 Other materials
Until now, we have considered systems where $J_{1}$-$J_{2}$ chains with a bipartite structure transition to CAFO under AFM interchain coupling. However, a simple bipartite structure is not that necessary. For instance, the Q1D mineral antlerite Cu3SO4(OH)4. The magnetic model of this system can be represented as three-leg zigzag ladders, where the central leg hosts competing AFM exchanges, and each side leg features alternating NN FM exchanges $J_{1}$, $J_{1}^{\prime}$ and a weaker NNN AFM exchange $J_{2}$. These three chains are coupled by zigzag-shaped AFM interchain couplings. The side chains are dimerised FM-AFM $J_{1}$-$J_{2}$ chains, with parameters estimated as $J_{1}=-26$, $J_{1}^{\prime}=-11$, and $J_{2}=-6$ (in meV) by GGA+U approximation [73]. Without interchain coupling, the side chains would be in a spiral state [74, 75]. However, the experimentally observed low-temperature magnetic structure shows that the side legs are ferromagnetically polarised in opposite directions, while the central chain is in an AFM order. This suggests that the effect of interchain coupling causes the side chains to polarize ferromagnetically.
@@ 5 Conclusion
In this study, we have demonstrated a novel mechanism for the stabilisation of MBS in Q1D cuprates through interchain interactions. Unlike conventional MBS that require an external magnetic field, our results show that AFM interchain coupling can induce MBS even in the absence of an external field. This finding is significant as it provides a new perspective on the interplay between interchain and intrachain interactions in low-dimensional magnetic systems.
Our numerical simulations using the DMRG method confirm that the AFM interchain coupling can act similarly to an internal magnetic field, leading to the formation of CAFO and the subsequent stabilisation of MBS. This is particularly intriguing because the AFM interchain coupling has a dual role; while it destabilizes MBS at high fields, it stabilizes them at zero field.
Notably, we found that MBS can be stabilised in Li2CuO2 under ambient pressure and zero magnetic field, highlighting the robustness of this mechanism. However, this aspect has not been experimentally investigated, as it has not been viewed from this perspective before. Additionally, we explored the conditions under which this phenomenon occurs in other materials such as Ca2Y2Cu5O10, LiCuSbO4, and PbCuSO4(OH)2. Our findings suggest that the application of pressure or chemical substitution could further enhance the interchain coupling in these materials, making it feasible to experimentally observe the MBS predicted by our theoretical models.
@@ Methods
We employ the DMRG technique [76] to study the ground-state properties. For single-chain calculations, we study periodic chains with lengths up to $l_{x}=192$ keeping $\chi=800$ density-matrix eigenstates for renormalization procedure. The discarded weight is negligible. When the magnon binding energy is calculated, it is better to use periodic chains to avoid picking up low excitations that may occur near the edges of open chains. For two-chain calculations, we study open ladders with lengths up to $l_{x}=280$ keeping $\chi=3600$. The largest discarded weight is $w_{\rm d}\sim 10^{-11}$. To explicitly fix the direction of magnetisation, which corresponds to the direction in which spin-rotation symmetry is broken in each chain, magnetic fields are applied to the sites at both ends of an open ladder. The direction of the magnetic fields is parallel to the z-axis, with fields at both ends of a single chain aligned in the same direction, and those on adjacent chains in opposite directions. However, the magnitude of magnetisation extrapolated to the thermodynamic limit is independent of the values of these edge fields. For $12\times 8$-chain calculations, we keep $\chi=1200$ leading to $w_{\rm d}\sim 10^{-10}$. In this case, we focus on the collinear antiferromagnetic ordered state, where the quantum fluctuations are significantly suppressed, the DMRG accuracy is very good. Similar to the case of the open ladder, magnetic fields are applied to both ends of the system.
@@
We employ the dynamical DMRG (DDMRG) technique [77] to calculate the magnon excitation spectrum (MES), defined as We calculate the spin excitation spectrum
@@ Acknowledgments
We thank Ulrike Nitzsche for technical assistance. We thank S.D.Wilson and R.O.Kuzian for fruitful discussion. This project is funded by the German Research Foundation (DFG) via the projects A05 of the Collaborative Research Center SFB 1143 (project-id 247310070). We kindly acknowledge support by the (Polish) National Science Centre (NCN, Poland) under Project No. 2021/40/C/ST3/00177.
Authors’ contributions C.E.A. and S.N. performed numerical calculations and analysed the results. S.N. designed the study and drafted the manuscript in discussion with C.E.A. and S.-L.D.. All authors approved the final manuscript.
@@
[…]
Data availability The data analyzed in the current study are available at https://doi.org/10.5281/zenodo.14499485